# MEGS2: Memory-Efficient Gaussian Splatting via Spherical Gaussians and Unified Pruning

### v1

@@ Abstract
3D Gaussian Splatting (3DGS) has emerged as a dominant novel-view synthesis technique, but its high memory consumption severely limits its applicability on edge devices. A growing number of 3DGS compression methods have been proposed to make 3DGS more efficient, yet most only focus on storage compression and fail to address the critical bottleneck of rendering memory. To address this problem, we introduce MEGS2, a novel memory-efficient framework that tackles this challenge by jointly optimizing two key factors: the total primitive number and the parameters per primitive, achieving unprecedented memory compression. Specifically, we replace the memory-intensive spherical harmonics with lightweight arbitrarily-oriented spherical Gaussian lobes as our color representations. More importantly, we propose a unified soft pruning framework that models primitive-number and lobe-number pruning as a single constrained optimization problem. Experiments show that MEGS2 achieves a 50 static VRAM reduction and a 40% rendering VRAM reduction compared to existing methods, while maintaining comparable rendering quality.
@@ 1 Introduction
Although 3D Gaussian Splatting (3DGS) (Kerbl et al., 2023) has been rapidly replacing implicit neural radiance fields (NeRF) (Mildenhall et al., 2020) as the dominant paradigm in neural rendering—thanks to its fast reconstruction, real-time performance, and high-quality outputs—supporting its application across devices with varying constraints has become increasingly important. As a result, compression of 3DGS (Bagdasarian et al., 2025) is gaining significantly more attention, with the goal of enabling real-world use cases on edge devices, such as mobile 3D scanning and previewing, virtual try-on, and real-time rendering in video games.
Nevertheless, existing compression methods only focus on storage compression rather than memory compression. While the former storage compression speeds up the one-time data transfer of 3DGS files, the rendering memory dictates whether the 3DGS rendering process can run smoothly on edge devices. This limits the applicability of 3DGS applications to low-end devices like cell phones. For example, some methods based on neural compression, vector quantization, and hash grid compression (Chen et al., 2024; Girish et al., 2024; Lee et al., 2024) achieve high storage compression rates. However, these methods require decoding the full Gaussian parameters from a compressed state before rendering, even resulting in a larger rendering memory than the 3DGS methods without compression. Another branch of works, i.e., primitive pruning methods (Zhang et al., 2025b; Fang & Wang, 2024), are effective at reducing both storage and rendering memory due to the reduced primitives. However, these pruning methods still need to retain a substantial number of 3D Gaussians for a reasonable rendering quality, which still consumes a large rendering memory. How to further reduce the rendering memory of 3DGS-based rendering methods remains an open question.
@@ 3 Method
Overview Our pipeline (Fig.2) begins with standard 3DGS initialization and densification. To reduce the parameter count, we introduce a novel color representation for 3DGS by fully replacing spherical harmonics (SH) with spherical Gaussians (SG) (Section 3.2). The core of our approach is a unified soft pruning framework (Section 3.3). This framework treats VRAM compression as a constrained optimization problem, adjusting primitive opacities and lobe sharpness progressively. After this process, a final post-processing step removes near-invalid primitives and low-sharpness lobes to yield a highly compact set of Gaussian primitives that enables rendering with a significantly reduced VRAM footprint.
@@ 3.1 Preliminaries for 3D Gaussian Splatting
Modeling 3D Gaussian Splatting (3DGS) (Kerbl et al., 2023) represents a scene using a collection of point-based 3D continuous Gaussian primitives. Each primitive is defined by its center position , a positive definite covariance matrix , a global opacity , and a view-dependent color model. The covariance matrix is structured as to ensure its positive definiteness, where is a rotation matrix and is a scaling matrix. The view-dependent color is abstractly defined as a function
which maps a viewing direction from the unit sphere to an RGB color in .
Render Pipeline To render an image from a specific viewpoint, each primitive is projected onto the 2D image plane. The 2D projected covariance matrix is computed via , where represents the view transformation and is the Jacobian of the affine approximation of the projective transformation. The final pixel color is then calculated by blending all projected primitives that contribute to that pixel in a front-to-back order. The blending process is defined as:
where is the view-dependent color and is the rendering opacity of the -th primitive. This rendering opacity is determined by the primitive’s global opacity and the value of its projected 2D Gaussian at the pixel location. Specifically, the rendering opacity is calculated as:
where is the pixel coordinate, is the 2D projected mean, and is the 2D projected covariance matrix. The view-dependent color is a function of the viewing direction , commonly modeled using Spherical Harmonics (SH) coefficients.
@@ 3.2 Spherical Gaussians
Choice of arbitrarily-oriented SG lobes It is particularly notable that the lobe axes of different SG are not constrained to be orthogonal, nor is any regularization term introduced to enforce orthogonality. Instead, each SG lobe is allowed to have an arbitrary direction. This flexibility grants the SG model a higher degree of freedom, leading to greater representation capability compared to models with fixed orthogonal axes. As shown in Figure 3, SG-Splatting (Wang et al., 2024b) with fixed orthogonal axes demonstrated significant rendering performance degradation, further underscoring the importance of supporting arbitrarily oriented SG lobes.
@@ 3.3.1 Problem Formulation
We unify primitive count pruning and spherical lobe pruning into a memory-constrained optimization problem. The norm of the opacity vector represents the number of active primitives, as Gaussian primitives with zero opacity make no contribution to rendering. Similarly, the norm of the sharpness vector denotes the number of active spherical lobes, since lobes with zero sharpness exhibit no view-dependent effect and their color can be directly added to the diffuse term . Building upon GaussianSpa (Zhang et al., 2025b)’s sparsification framework and insights from our analysis, we extend the pruning objective from primitive count optimization to total memory budget control, formalized as:
@@ 3.3.2 Memory-constrained Optimization
Here, and denote the corresponding Lagrange multipliers for the constraints.
The specific algorithm flow can be found in Algorithm 1.
@@
For opacity projection, Zhang et al. (2025b) has established multiple proximal operators that are selected based on scene characteristics. Our approach reuses these operators.
@@ 3.3.3 Post-Processing Procedure
Solving this optimization problem yields an exact compensation term that preserves the energy of the removed lobe:
@@
Finally, after removing some Gaussian primitives and spherical lobes, we continue to fine-tune for a small number of steps to recover the rendering quality.
@@ 4.1 Experimental Settings
Implementation Details Our experiments are conducted on a single NVIDIA RTX 3090 GPU (24GB). For a fair comparison, we adopt the same hyperparameters and optimizer used by vanilla 3DGS on all datasets. Regarding the training strategy, we employ the same densification, pruning, and fine-tuning procedures as GaussianSpa, conducting densification training, sparse training, and fine-tuning for the same number of steps. To ensure the accuracy of our VRAM measurements, we reproduce 3DGS, GaussianSpa, EAGLES, CompactGaussian, MesonGS, and Reduced3DGS for direct VRAM measurement. For other pruning-based methods, we utilize both the rendering quality metrics and primitive counts reported in the GaussianSpa paper. We then estimate their corresponding VRAM consumption based on our own VRAM measurements for 3DGS and GaussianSpa.
@@ 4.2 Main Results and Analysis
The quantitative results are summarized in Table 1, comparing our approach with several existing lightweight 3DGS methods. Overall, our method achieves comparable or slightly superior rendering quality to the current state-of-the-art lightweight 3DGS methods, while demonstrating significantly lower VRAM overhead across all existing methods.
Compared to vanilla 3DGS, our method achieves a 0.4 dB PSNR improvement on the DeepBlending dataset and a 0.012 LPIPS reduction on the Tanks & Temples dataset. Furthermore, it achieves more than an 8 compression rate for static VRAM and nearly a 6 compression rate for rendering VRAM across all datasets.
Even when compared to the current state-of-the-art lightweight 3DGS method, GaussianSpa (Zhang et al., 2025b), we still achieve a nearly 2 compression rate for static VRAM on the Mip-NeRF360 dataset and reduce rendering VRAM by approximately 40% with comparable rendering quality.
Analysis on Existing Compression Methods It is noteworthy that methods based on vector quantization, neural network compression, or hash grid compression (CompactGaussian (Lee et al., 2024), EAGLES (Girish et al., 2024), MesonGS (Xie et al., 2024), despite achieving high storage compression rates, lack advantages in terms of static and rendering VRAM. We summarize the average parameter costs for the color model of these methods in Table 2. This is because they still cannot render directly from their compressed state in the latent space or codebook space; they must first decode the full Gaussian parameters. This decoding and dequantization overhead results in both their static and rendering VRAM being even higher than that of vanilla 3DGS methods with a similar number of primitives. Similarly, we compare our method against HAC++ (Chen et al., 2025), a state-of-the-art storage compression algorithm only for anchor-based 3DGS on the Mip-NeRF360 dataset. As shown in Table 3, HAC++ does not significantly reduce VRAM compared to its baseline, Scaffold-GS (Lu et al., 2023), because its compressed representation must be fully decompressed for rendering. In contrast, our method demonstrates a clear advantage in metrics closer to human perception (SSIM and LPIPS) while achieving approximately 50-60% lower VRAM overhead and a 1.5-1.7x rendering speedup.
Analysis on Pruning-based Methods Pruning-based methods (Zhang et al., 2025b; Fang & Wang, 2024; Zhang et al., 2024), while effectively reducing both static and rendering VRAM, have reached a point where it is difficult to further reduce the number of Gaussian primitives without compromising rendering quality, thus limiting the potential for higher VRAM compression rates. For methods that independently prune primitives and spherical harmonics, such as Reduced3DGS (Papantonakis et al., 2024), our advantage stems from two key factors: the superior pruning amenability of Spherical Gaussians and the more effective, unified soft pruning framework. A detailed discussion and experimental validation of this superiority are provided in Section 4.4.
@@ 4.3 Qualitative Results
Figure 4 presents a qualitative comparison of our method against baselines, including GaussianSpa and 3DGS on various scenes. In the Bicycle and Truck scenes, our method accurately recovers specular reflections on smooth surfaces and mirrors, details that other methods fail to fully capture. In the Bonsai scene, our approach faithfully captures high-contrast lighting in both contours and brightness. Furthermore, in the Playroom scene, where GaussianSpa struggles with the reconstruction of the wall corner, and 3DGS suffers from a loss of detail in other regions, our method delivers a cleaner and more complete reconstruction. Notably, our method achieves these high-quality results with a VRAM footprint of only 50-60% compared to GaussianSpa, demonstrating a substantial enhancement in memory efficiency. These visual results indicate that our Spherical Gaussian (SG) based representation has a stronger capability for fitting local view-dependent signals than Spherical Harmonics (SH), leading to sharper and more photorealistic images.

### v2

@@ Abstract
3D Gaussian Splatting (3DGS) has emerged as a dominant novel-view synthesis technique, but its high memory consumption severely limits its applicability on edge devices. A growing number of 3DGS compression methods have been proposed to make 3DGS more efficient, yet most only focus on storage compression and fail to address the critical bottleneck of rendering memory. To address this problem, we introduce MEGS2, a novel memory-efficient framework that tackles this challenge by jointly optimizing two key factors: the total primitive number and the parameters per primitive, achieving unprecedented memory compression. Specifically, we replace the memory-intensive spherical harmonics with lightweight, arbitrarily oriented spherical Gaussian lobes as our color representations. More importantly, we propose a unified soft pruning framework that models primitive-number and lobe-number pruning as a single constrained optimization problem. Experiments show that MEGS2 achieves a 50 static VRAM reduction and a 40% rendering VRAM reduction compared to existing methods, while maintaining comparable rendering quality. Project page: https://megs-2.github.io/
@@ 1 Introduction
Although 3D Gaussian Splatting (3DGS) (Kerbl et al., 2023) has been rapidly replacing implicit neural radiance fields (NeRF) (Mildenhall et al., 2020) as the dominant paradigm in neural rendering, thanks to its fast reconstruction, real-time performance, and high-quality outputs, supporting its application across devices with varying constraints has become increasingly important. As a result, compression of 3DGS (Bagdasarian et al., 2025) is gaining significantly more attention, with the goal of enabling real-world use cases on edge devices, such as mobile 3D scanning and previewing, virtual try-on, and real-time rendering in video games.
Nevertheless, existing compression methods only focus on storage compression rather than memory compression. While the former storage compression speeds up the one-time data transfer of 3DGS files, the rendering memory dictates whether the 3DGS rendering process can run smoothly on edge devices. This limits the applicability of 3DGS applications to low-end devices like cell phones. For example, some methods based on neural compression, vector quantization, and hash grid compression (Chen et al., 2024; Girish et al., 2024; Lee et al., 2024) achieve high storage compression rates. However, these methods require decoding the full Gaussian parameters from a compressed state before rendering, even resulting in a larger rendering memory than the 3DGS methods without compression. Another branch of works, i.e., primitive pruning methods (Zhang et al., 2025b; Fang & Wang, 2024), is effective at reducing both storage and rendering memory due to the reduced primitives. However, these pruning methods still need to retain a substantial number of 3D Gaussians for a reasonable rendering quality, which still consumes a large rendering memory. How to further reduce the rendering memory of 3DGS-based rendering methods remains an open question.
@@ 3.1 Preliminaries for 3D Gaussian Splatting
3D Gaussian Splatting (3DGS) (Kerbl et al., 2023) represents a scene using a set of 3D Gaussian primitives. Each primitive is defined by a center , a covariance matrix , a scalar opacity , and a view-dependent color model. The color is determined by a function , which maps a viewing direction to an RGB color. In the original work, this function is modeled using Spherical Harmonics (SH).
To render an image, these 3D primitives are projected onto the 2D image plane. The final pixel color is computed by alpha-blending the primitives that overlap with the pixel in front-to-back order, according to the equation: . Here, the rendering opacity is a product of the primitive’s scalar opacity and its 2D Gaussian value at the pixel location.
@@ 3.2 Spherical Gaussians
Choice of arbitrarily-oriented SG lobes It is particularly notable that the lobe axes of different SG are not constrained to be orthogonal, nor is any regularization term introduced to enforce orthogonality. Instead, each SG lobe is allowed to have an arbitrary direction. This flexibility grants the SG model a higher degree of freedom, leading to greater representation capability compared to models with fixed orthogonal axes. As shown in Figure 4 (in Appendix), SG-Splatting (Wang et al., 2024b) with fixed orthogonal axes demonstrated significant rendering performance degradation, further underscoring the importance of supporting arbitrarily oriented SG lobes.
@@ 3.3.1 Problem Formulation
We unify primitive count pruning and spherical lobe pruning into a memory-constrained optimization problem. The norm of the opacity vector represents the number of active primitives, as primitives with zero opacity do not contribute to rendering. Similarly, the norm of the sharpness vector denotes the number of active spherical lobes, since lobes with zero sharpness exhibit no view-dependent effect and their color can be directly added to the diffuse term . Building upon GaussianSpa (Zhang et al., 2025b)’s sparsification framework and our analysis, we extend the pruning objective from primitive count optimization to total memory budget control, formalized as:
@@ 3.3.2 Memory-constrained Optimization
Here, and denote the corresponding Lagrange multipliers for the constraints. The specific algorithm flow can be found in Algorithm 1, and the detailed derivation is in Appendix A.1.1.
@@
For opacity projection, Zhang et al. (2025b) has established multiple proximal operators that are selected based on scene characteristics. Our approach reuses these operators (See Appendix A.2).
@@ 3.3.3 Post-Processing Procedure
Solving equation 11 yields an exact compensation term that preserves the color of the removed lobe:
@@
Finally, after removing some primitives and spherical lobes, we continue to fine-tune for a small number of steps to recover the rendering quality. The detailed derivation is in Appendix A.1.2
@@ 4.1 Experimental Settings
Experimental Setup To ensure a fair comparison, we follow the core training procedures established by 3DGS (Kerbl et al., 2023) and GaussianSpa (Zhang et al., 2025b). We evaluate rendering quality using PSNR, SSIM, and LPIPS, and measure efficiency via static and rendering VRAM consumption. For a detailed description of our specific implementation, as well as the definitions and measurement procedures for both static and rendering VRAM, please refer to Appendix A.2.
@@ 4.2 Main Results and Analysis
The quantitative results are summarized in Table 1, comparing our approach with several existing lightweight 3DGS methods. Overall, our method achieves comparable or slightly superior rendering quality to the current state-of-the-art lightweight 3DGS methods, while demonstrating significantly lower VRAM overhead across all existing methods. Compared to vanilla 3DGS, our method achieves a 0.4 dB PSNR improvement on the DeepBlending dataset and a 0.012 LPIPS reduction on the Tanks & Temples dataset. Furthermore, it achieves more than an 8 compression rate for static VRAM and nearly a 6 compression rate for rendering VRAM across all datasets. Even when compared to the current state-of-the-art lightweight 3DGS method, GaussianSpa (Zhang et al., 2025b), we still achieve a nearly 2 compression rate for static VRAM on the Mip-NeRF360 dataset and reduce rendering VRAM by approximately 40% with comparable rendering quality.
Analysis on Existing Compression Methods Many recent methods achieve high storage compression but fail to reduce VRAM. Techniques based on hash grid compression, vector quantization or neural networks (e.g., CompactGaussian (Lee et al., 2024), EAGLES (Girish et al., 2024)) must first decompress their parameters into a renderable state. This requirement, coupled with the decoding process itself, results in a VRAM footprint that can exceed even vanilla 3DGS (Appendix, Tab. 3). Similarly, HAC++ (Chen et al., 2025), a SOTA method designed specifically for anchor-based 3DGS, is not a general solution and offers minimal VRAM reduction for the same reason. In contrast, our method achieves superior perceptual quality (SSIM/LPIPS) with 50-60% less VRAM and a 1.5-1.7x rendering speedup.
Analysis on Pruning-based Methods Pruning-based methods (Zhang et al., 2025b; Fang & Wang, 2024; Zhang et al., 2024) effectively reduce VRAM but face a bottleneck, as aggressively reducing primitive count degrades quality. Our method overcomes this by not only pruning primitives but also reducing per-primitive costs. When compared to methods that also compress color, such as Reduced3DGS (Papantonakis et al., 2024) which prunes SH coefficients, our advantage is twofold: the superior amenability of our Spherical Gaussians to pruning, and our more effective unified soft pruning framework, as validated in Section 4.4.
@@ 4.3 Qualitative Results
Figure 3 presents a qualitative comparison of MEGS2 against baselines, including GaussianSpa and 3DGS on various scenes. In Bicycle and Truck, our method accurately recovers specular reflections on smooth surfaces and mirrors, details that other methods fail to fully capture. In Bonsai, MEGS2 faithfully captures high-contrast lighting in both contours and brightness. Furthermore, MEGS2 delivers a cleaner and more complete reconstruction in Playroom. Notably, MEGS2 achieves these high-quality results with a VRAM footprint of only 50-60% compared to GaussianSpa, demonstrating a substantial enhancement in memory efficiency. These visual results indicate that our Spherical Gaussian based representation has a stronger capability for fitting local view-dependent signals than Spherical Harmonics, leading to sharper and more photorealistic images.
4.4 Ablation Studies
In this section, we conduct a series of ablation studies to validate the effectiveness of our proposed designs and answer two key questions: 1. Is the replacement of spherical harmonics with spherical Gaussians necessary? And why not use the theoretically lighter-weight orthogonal-axis spherical variants? (Section 4.4.1) 2. Is our proposed unified soft pruning framework effective? And does it offer a significant advantage over simply combining existing pruning methods? (Section 4.4.2)
4.4.1 Ablation on Color Modeling
To validate our color model, we replace the spherical harmonics (SH) in 3DGS with our arbitrarily-oriented spherical Gaussians (SG) while keeping training settings identical. Our SG model achieves a superior quality-VRAM trade-off (see Appendix, Fig. 4). Furthermore, we show that the fixed-axis SG from SG-Splatting (Wang et al., 2024b) struggles to capture complex view-dependent effects, leading to a sharp quality drop (about 0.6 dB PSNR). This rigidity also makes it incompatible with our lobe-pruning strategy.
| Method | PSNR | SSIM | LPIPS | VRAM |
| Spa+Redu. | 26.05 | 0.776 | 0.280 | 402 |
| Spa(SH SG) | 27.01 | 0.808 | 0.230 | 339 |
| soft hard | 27.23 | 0.814 | 0.228 | 288 |
| unified seq. | 27.33 | 0.818 | 0.222 | 328 |
| w/o color comp. | 27.46 | 0.822 | 0.213 | 265 |
| full model | 27.54 | 0.824 | 0.209 | 265 |
4.4.2 Ablation on Pruning Strategies
We evaluate our unified soft pruning framework against several ablations and baselines, with results in Table 2. First, replacing our soft pruning with a hard-pruning variant (soft → hard) that only optimizes opacity results in a significant performance drop. Second, performing primitive and lobe pruning sequentially (unified → sequential) is inferior to our joint optimization, confirming the advantage of a unified approach. We also found that removing our color compensation harms performance (w/o color comp.). Finally, we test two naive baselines. Combining GaussianSpa with Reduced3DGS leads to severe quality degradation, showing that existing SH pruning metrics fail on fewer primitives. Simply replacing SH with SG in GaussianSpa is also outperformed by our method. This demonstrates that our pruning strategy not only reduces memory but also acts as a regularizer, improving rendering quality.
5 Conclusion
We introduced MEGS2, a framework that addresses the 3D Gaussian Splatting memory bottleneck by jointly optimizing primitive count and per-primitive parameters. Our approach combines a lightweight spherical Gaussian color representation with a unified soft pruning method to achieve state-of-the-art memory compression. MEGS² shifts the focus of 3DGS compression towards rendering VRAM efficiency, paving the way for high-quality rendering on edge devices.
Limitation Our method focuses on compressing static memory for broad, renderer-agnostic applicability. We leave the optimization of implementation-specific dynamic VRAM for future work, and note that the model’s performance on highly complex highlights warrants further investigation.
Ethics statement
Our work presents no direct ethical concerns. The primary application of our method is for novel view synthesis from captured data.
Reproducibility statement
To ensure reproducibility: (1) While not included with this submission, our full project (including the WebGL renderer, training, and evaluation scripts) will be released on GitHub upon publication. (2) All experimental details, including our VRAM measurement protocol, are provided in Appendix A.2. (3) Our data and preprocessing follow the official 3DGS implementation. (4) Derivations and proofs for our core algorithms are in Appendix A.1.1 and A.1.2.
Appendix A Appendix
A.1 Algorithmic Derivation
A.1.1 Derivation for Memory-constrained optimizing
This appendix provides the detailed derivation of Algorithm 1 presented in the main text. The algorithm solves the memory-constrained optimization problem with norm constraints using an ADMM-inspired approach.
Problem Formulation Consider the constrained optimization problem:
| s.t. |
where is the loss function, denotes the norm (count of non-zero elements), and represents the memory budget.
Augmented Lagrangian Formulation Introduce proxy variables , to reformulate the problem:
| s.t. | ||||
The augmented Lagrangian function is constructed as:
where and are Lagrange multipliers, and is the penalty parameter.
ADMM Alternating Minimization The ADMM framework decomposes the optimization into three subproblems:
Primal Variables Update With proxy variables and multipliers fixed, update the primal variables:
Using gradient descent approximation:
where .
Proxy Variables Update With primal variables and multipliers fixed, update the proxy variables:
where is the indicator function enforcing the sparsity constraint:
The solution is given by the proximal operator:
Dual Variables Update Update the Lagrange multipliers:
A.1.2 Derivation for Color Compensation
Simplifying the objective function:
Expanding the squared term:
Taking derivative with respect to and setting to zero:
Solving for the optimal :
Evaluating the spherical Gaussian integral. Without loss of generality, align with the z-axis:
Substituting back:
The diffuse term is updated as:
A.2 Experimental Details
Implementation Details Our experiments are conducted on a single NVIDIA RTX 3090 GPU (24GB). For a fair comparison, we adopt the same hyperparameters and optimizer used by vanilla 3DGS (Kerbl et al., 2023) on all datasets. Regarding the training strategy, we employ the same densification, pruning, and fine-tuning procedures as GaussianSpa, conducting densification training, sparse training, and fine-tuning for the same number of steps. To ensure the accuracy of our VRAM measurements, we reproduce 3DGS, GaussianSpa (Zhang et al., 2025b), EAGLES (Girish et al., 2024), CompactGaussian (Lee et al., 2024), MesonGS (Xie et al., 2024), and Reduced3DGS (Papantonakis et al., 2024) for direct VRAM measurement. For other pruning-based methods, we utilize both the rendering quality metrics and primitive counts reported in the GaussianSpa paper. We then estimate their corresponding VRAM consumption based on our own VRAM measurements for 3DGS and GaussianSpa. EAGLES cannot be trained on a single NVIDIA RTX 3090 GPU with 24GB on some scenes in Mip-NeRF360, so we do not report VRAM consumption.
Proximal Operators Implementation As mentioned in the main text, for opacity projection, we reuse the two proximal operators from GaussianSpa (Zhang et al., 2025b). These operators select primitives to preserve based on different importance criteria. The first is a Magnitude-Based Selection, which sorts the input vector and preserves the elements with the highest magnitudes. The second operator employs an Importance-Based Selection, adopting the importance score metric from MiniSplatting (Fang & Wang, 2024). The importance score for the -th Gaussian is the sum of its blending weights across all intersecting camera rays seen during training:
where the blending weight is the product of the Gaussian’s opacity , its projected 2D value , and the accumulated transmittance along the -th ray. Here, is the total number of intersected rays. The operator then preserves the primitives with the highest importance scores.
A.3 More Quantitive Results